OpenStudy (anonymous):

4 5/6+9+12 5/6-4-3 5/6 Help! I wanna see if me or my teacher is right I got 18 5/6 and she got 15 5/6 Every time i do it i get 18 5/6

TheSmartOne (thesmartone):

Could you write that using the equation button>

OpenStudy (anonymous):

sure :)

TheSmartOne (thesmartone):

(:

OpenStudy (anonymous):

\[4 \frac{ 5 }{ 6}+9+12\frac{ 5 }{ 6 }-4-3\frac{ 5 }{ 6 }=?\]

TheSmartOne (thesmartone):

\[4 \frac{ 5 }{ 6}+9+12\frac{ 5 }{ 6 }-4-3\frac{ 5 }{ 6 }=?\] \[\frac{ 29 }{ 6}+9+\frac{ 77 }{ 6 }-4-\frac{ 23 }{ 6 }=?\]

OpenStudy (anonymous):

Oh i didnt make them improper fractions lol

TheSmartOne (thesmartone):

\[ 9-4 + \frac{29+77-23}{6}\]

OpenStudy (anonymous):

18.83333333333333

TheSmartOne (thesmartone):

\[ 9-4 + \frac{83}{6}\]

TheSmartOne (thesmartone):

\[ 5+ \frac{83}{6}\]

TheSmartOne (thesmartone):

\[ 5+ 13\frac{5}{6}\]

TheSmartOne (thesmartone):

@angel12310 Tell your teacher to retire.

OpenStudy (anonymous):

So im right!

TheSmartOne (thesmartone):

I think she's getting old. ;P

OpenStudy (anonymous):

She's not old lol!

OpenStudy (anonymous):

Thanks for the help!

TheSmartOne (thesmartone):

No problem. \(\ddot\smile\)

OpenStudy (texaschic101):

4 5/6 + 9 + 12 5/6 - 4 - 3 5/6 29/6 + 9 + 77/6 - 4 - 23/6 29/6 + 54/6 + 77/6 - 24/6 - 23/6 160/6 - 47/6 = 113/6 = 18 5/6 or 18.833

OpenStudy (anonymous):

Thanks @texaschic101

OpenStudy (texaschic101):

no problem :)

OpenStudy (anonymous):

I wish i can give u a medal 2!